OpenStudy (lilai3):

Please help! Will give a medal! You will not regret it! LOOOL But Please Help!!!!! In a month with 30 days, what percent of the dates are prime numbers?

OpenStudy (anonymous):

how many numbers are prime from 1 to 30?

OpenStudy (lilai3):

i think 12: 1 2 3 5 7 9 11 13 17 19 23 29

OpenStudy (lilai3):

12/30=4/10=0.40=40%

OpenStudy (anonymous):

to find a percentage divide the total by the number of primes, then multiply by 100 100(total days/numberofprimes)

OpenStudy (anonymous):

yeah your right im wrong :)

OpenStudy (lilai3):

oh okay thank you bye thank you for help. and oh yes. your medal! ;)

OpenStudy (lilai3):

thanks again. have a good day...or night..it's night here in my hometown... lols bye anyways